# Teleportation Fidelity of Binary Tree Quantum Repeater Networks

### v2

@@ Abstract
The idea of average of maximum teleportation fidelities was introduced in [1] to measure the capability of the network to act as a resource for distributed teleportation between any pair of nodes. Binary tree network, being a subclass of Cayley tree network, is a significant topological structure used for information transfer in a hierarchical sense. In this article, we consider four types of binary tree repeater networks (directed and undirected, asymmetric and symmetric) and obtain the analytical expressions of the average of the maximum teleportation fidelities for each of these binary tree networks. Our work investigates the role of directionality and symmetry in this measure. Like [1], here also we have used simple Werner state-based models and are able to identify the parameter ranges for which these networks can show quantum advantage. We also explore the role of maximally entangled states in the network to enhance the quantum advantage. We provide a detailed examination of the large-scale behavior of these networks, obtaining the limiting value of the average maximum teleportation fidelity as the number of nodes, $N$, approaches infinity. To model realistic conditions, we also analyze scenarios where the Werner state parameters for each link are different and chosen from a uniform distribution. Our findings reveal that the directed symmetric binary tree represents the most advantageous topology for quantum teleportation within this context. The work makes a significant advancement in the process of identifying resourceful tree networks for distributed quantum teleportation. We also contribute a methodology for the analytical calculation of pathlengths in all considered graph types, which can pave the way for future studies on more general Cayley trees with arbitrary degrees.
I Introduction
@@
In the last decade, we have seen a steady rise in the research and implementation of quantum networks with the vision of creating a full-stack quantum internet [2, 3, 4, 5]. Quantum networking also becomes crucial in distributed quantum computing, where separated quantum computers can work on a similar algorithm [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Not only computing, but quantum networks can become a test bed for various information processing tasks with many more parties and distributed scenarios. In that spirit, we have seen network topology to play a significant role and value-addition in choosing a network best suited for a particular task [1]. Unlike classical network [18, 19, 20, 21, 22, 23, 24, 25, 26, 27], one of the unexplored areas is the large-scale quantum networks. For a practical purpose, quantum networks can be deployed either through satellite-based free space technology [28, 29] or large ground-based repeater technology [30, 31, 32, 17]. In an entanglement-based repeater network, entanglement, being the key ingredient, have a crucial role in various communication tasks like sending quantum information, classical information, key generation [33, 34, 35] and secret sharing [36, 37, 38, 39, 40, 41, 42, 43, 34, 44, 45, 46, 47, 48, 49, 50]. In such networks the distribution of entanglement (both bipartite [32, 17] and multipartite [51]) becomes a crucial factor in spite of limitations of achieving relevant quantum advantages [52, 53, 12]. In this article, we consider a ground-based repeater network specially curated for the purpose of teleportation.
@@
In a repeater-based quantum network scenario, establishment of entanglement between two distant nodes [32, 59, 17] happens through the process of entanglement swapping in the repeater stations [60]. Given the fidelity of intermediate states, it becomes interesting to find out the teleportation fidelity of the entangled state established between the source and target as a result of swapping at the repeater station (Fig. 1). In a quantum repeater network made up of Werner states (parametrized by a set of values $p_{i}$), the maximum teleportation fidelity through a particular path (P) connecting the source (S) and the target (T) can be calculated $F^{\text{tel}}$ as $F^{\rm max}_{{\rm ST},\mathcal{P}}(\rho_{\rm wer})=\left(1+\prod_{i\in\mathcal{P}}p_{i}\right)/2$ [60, 61]. In fact, this is the teleportation fidelity of the final entangled state between S and T created by swapping through that path. In general, in a full-scale quantum network, there can be multiple paths from a source to a target, and hence multiple teleportation fidelities between S and T based on different paths can exist. However, we take the maximum of all the teleportation fidelities when we fix the source and target [1]. To quantify the resourcefulness of a network globally in the context of teleportation, a measure called average of maximum teleportation fidelity ($F^{\text{tel}}_{\text{avg}}$) achieved by all paths was introduced in a recent work [1]. This has given us a global metric ($F^{\text{tel}}_{\text{avg}}$) to estimate the capability of a network in carrying out teleportation over any source and target, and eventually ranking them. In that article [1], some of us have investigated how teleportation fidelity (on an average scale) varies in both loop and loopless networks. We have examined tree sequences that range from chains with all nodes having degree two (apart from the extreme two nodes) to stars, where the hub has the most links. Other trees, like branching trees, that can be created by gradually rearranging nodes from the chain, have also been investigated [1]. It’s been demonstrated that the network can achieve the quantum average fidelity with a lower average distance.
@@
Motivated by this newly introduced measure $F_{\text{avg}}^{\text{tel}}$ in undirected simple trees, we focus here on a class of directed and undirected binary trees, a subclass of the Cayley tree. A Cayley tree is an acyclic, connected graph in which each node has up to $k$ children. The binary trees are a specific subcategory of Cayley tree. A left child and a right child are the maximum number of offspring that each node in the binary tree, a hierarchical tree structure, can have. Note that binary trees are used in computer science for designing algorithms and data structures [62, 63]. A lot of studies have been conducted to investigate the role of (un)directed trees on percolation [64, 65, 66, 67, 68, 69], synchronization [70, 71], ecological stability, and in the immune network model  [72, 73]. Investigating the average of maximum teleportation fidelity in directed and undirected binary trees is necessary to understand the process of quantum information transfer in a hierarchical network structure and, hence, will be the focus of the current work. Specifically, both symmetric and asymmetric binary trees will be used to study the $F_{\text{avg}}^{\text{tel}}$ in this article. In the symmetric situation, each node has two children (left and right) at each hop/level. The right child will not be taken into consideration if each level has an asymmetric instance. The current work unfolds two aspects: (i) How does symmetry in a binary graph shape the global teleportation fidelity, and (ii) how does directionality in links influence the global behavior?
@@
In this article, we consider four types of binary trees: directed asymmetric binary tree (DABT), directed symmetric binary tree (DSBT), undirected asymmetric binary tree (UABT) and undirected symmetric binary tree (USBT). We analytically and numerically find out the average of maximum teleportation fidelities for all these binary tree topologies (See Section II.3, Fig. 2). Our network models are based on edges made up of Werner states parametrized by the parameter $p$ uniformly over the network. We also obtain the range of $p$ for which each networks show quantum advantage with varying network size $N(d)$, (parametrized by depth $d$). In that process, we also find out the limiting value of the average of the maximum teleportation fidelity for each of these networks for a large number of $N(d)$. Given that each Werner state is useful for teleportation for $p>1/3$, we find the lower bound of the average of the maximum teleportation fidelity for each network and capture its variation with increasing $N(d)$.
@@
In Section II, we give a brief introduction to existing concepts like entanglement swapping, average of maximum teleportation fidelity and binary tree network. In Section III, we obtain the average of maximum teleportation fidelity for each of these four binary networks. In Section IV, we analyze the quantum advantage for each of these networks and also find out the lower bounds of the average of maximum teleportation fidelities when each link is useful for teleportation. In Section V, we have calculated the average of maximum fidelity when the input state parameter of the Werner states is different for each edge of the corresponding binary tree network (the input parameters are chosen from a uniform distribution). Finally, in Section VI, we have concluded by summarizing the findings of this work.
@@
Figure 1: Entanglement Swapping: (a) Alice and Charlie are not connected, but entangled pairs are shared between Alice–Bob and Bob–Charlie. (b) Bob’s measurement creates entanglement between Alice and Charlie. (c) Sequential protocol to establish entanglement between $N+2$ nodes using NN swaps.
II Preliminaries
@@
II.1 Entanglement Swapping and Quantum Repeater
Entanglement swapping is the process by which you can teleport a quantum entanglement from one location to another, thus by connecting two nodes which are not initially entangled with each other. The simplest scheme of entanglement swapping involves three parties, Alice, Bob and Charlie. Here Alice and Charlie shares Bell state $|\Phi^{+}\rangle_{AB_{1}}$ and $|\Phi^{+}\rangle_{B_{2}C}$ with Bob. Now, if Bob carries out two two-qubit Bell-state measurements on his qubits $B_{1}$ and $B_{2}$, Alice and Charlie will end up with two two-qubit states depending upon the measurement outcomes (See Fig. 1) at Bob’s location,
@@ II.1 Entanglement Swapping and Quantum Repeater
as a resource between Alice-Bob and Bob-Charlie, respectively. After first swapping the state between Alice and Charlie will become [60, 61],
@@
If we extend this idea to a $(N+2)$ node scenario, to establish the connection between 11st and the $(N+2)$th node , then there should be $N$ such swappings (See Fig. 1(c)). The resultant state will become,
@@
II.2 Teleportation Fidelity and Average of Maximum Teleportation Fidelity
@@
II.3 Binary Tree Network
A binary tree is one of the hierarchical tree data structures, with each node having at most two child nodes. This form of tree is used in search engines, web page organization and indexing, network data routing, decision analysis, and [62, 63]. Based on the structure, there are two types of binary trees: asymmetric and symmetric. Similarly, based on the directionality of the edges, we can classify binary trees into two types: directed and undirected. Here in this article, we consider four different types of binary trees: directed asymmetric and symmetric binary tree (DABT, DSBT) and undirected asymmetric and symmetric binary tree (UABT, USBT).
@@ II.3 Binary Tree Network
In Figs. 2(a), 2(b), 2(c), and 2(d), we showed four different cases of binary tree network i.e. directed asymmetric binary tree (DABT), directed symmetric binary tree (DSBT), undirected asymmetric binary tree (UABT), and undirected symmetric binary tree (USBT) respectively. The root node (depicted in skyblue colour) of the tree is node number 1 in each of the cases. In the case of the asymmetric trees (DABT and UABT), the leaf nodes (depicted in green colour) are 3, 5, 6, and 7, whereas, the leaf nodes are 8, 9, 10, 11, 12, 13, 141 and 15 in the case of symmetric trees (DSBT and USBT). The depth of each tree is $d=3$ in Figs. 2(a), (b), (c), and (d). The total number of nodes in DABT, DSBT, UABT and USBT is 7, 15, 7, and 15, respectively.
@@
Figure 2: From Left to Right: We have DABT, DSBT, UABT and USBT with 7, 15, 7 and 15 nodes, respectively. The depth of each binary tree is $d$ = 3.
III Average Teleportation Fidelity in Binary Tree Network
In this section, we consider a repeater-based entangled network of $N$ nodes arranged in a binary tree structure. All the edges of our network are made up of Werner states given in Eq. (2) parameterized by the parameter $p$. In this section, we consider four cases of the binary tree network: 1) directed asymmetric binary tree (DABT), 2) directed symmetric binary tree (DSBT), 3) undirected asymmetric binary tree (UABT), 4) undirected symmetric binary tree (USBT) and obtain the analytical expression for average teleportation fidelity for each of these networks.
III.1 Directed Binary Tree
In this subsection, we consider two types of directed binary trees. These include both asymmetric and symmetric cases.
@@ III.1 Directed Binary Tree
$\left.F^{\text{tel}}_{\text{avg}}(d,p)\right|_{\text{DABT}}$ $=$ $\frac{\sum\limits_{i=1}^{d}[2+2(d-i)]\mathcal{F}_{i}}{\sum\limits_{i=1}^{d}[2+2(d-i)]}$ (7)
$=$ $\frac{1}{2}+\frac{p[p(p^{d}-1)+d(1-p)]}{d(d+1)(1-p)^{2}}.$ (8)
@@
$\left.F^{\text{tel}}_{\text{avg}}(N,p)\right|_{\text{DABT}}=\frac{\sum\limits_{i=1}^{(N-1)/2}(N+1-2i)\mathcal{F}_{i}}{\sum\limits_{i=1}^{(N-1)/2}(N+1-2i)}.$ (9)
$=$ $\frac{1}{2}+\frac{2p[2p(p^{\frac{N-1}{2}}-1)+(N-1)(1-p)]}{(N-1)(N+1)(1-p)^{2}}.$ (10)
@@
$\left.F^{\text{tel}}_{\text{avg}}(d,p)\right|_{\text{DSBT}}$ $=$ $\frac{\sum\limits_{i=1}^{d}[2^{d+1}-2^{i}]\mathcal{F}_{i}}{\sum\limits_{i=1}^{d}[2^{d+1}-2^{i}]}=\begin{cases}\displaystyle\frac{1}{2}+\frac{2^{d+1}-d-2}{2\left[(d-1)\cdot 2^{d+1}+2\right]},&\text{if }\leavevmode\nobreak\ p=\frac{1}{2},\\[12.0pt] \displaystyle\frac{1}{2}+\frac{p\left[(2^{d}-1)-p(2^{d+1}-1)+2^{d}p^{d+1}\right]}{2(1-p)(1-2p)\left[(d-1)\cdot 2^{d}+1\right]},&\text{otherwise},\end{cases}$ (11)
@@
$\left.F^{\text{tel}}_{\text{avg}}(N,p)\right|_{\text{DSBT}}=\frac{\sum\limits_{i=1}^{\log_{2}[(N+1)/2]}(N+1-2^{i})\mathcal{F}_{i}}{\sum\limits_{i=1}^{\log_{2}[(N+1)/2]}(N+1-2^{i})}=\begin{cases}\displaystyle\frac{1}{2}+\frac{N-\log_{2}(N+1)}{2\left[(N+1)\log_{2}(N+1)-2N\right]},&\text{if }\leavevmode\nobreak\ p=\frac{1}{2},\\[12.0pt] \displaystyle\frac{1}{2}+\frac{p[N(1-2p)+(N+1)p^{\log_{2}(N+1)}-1]}{4(1-p)(1-2p)\left[(N+1)\log_{2}(N+1)-2N\right]},&\text{otherwise}.\end{cases}$ (12)
A detailed derivation of the formula is provided in appendix A.1.
III.2 Undirected Binary Trees
Undirected Asymmetric Binary Trees (UABT) : Here we consider an undirected asymmetric binary tree network (UABT) where each link represents a Werner state, where each Werner state is parameterized by the classical mixing parameter $p$. It is shown in Fig. 2(c). In the UABT network, we start with $3$ nodes at level $h=1$, and after that we add two nodes at every level. So in that case $N$ will be odd numbers starting from $3$ (i.e. $N=3,5,7\cdots$).

### v3

@@ Abstract
Binary tree network, being a subclass of Cayley tree network, is a significant topological structure used for information transfer in a hierarchical sense. In this article, we consider four types of binary tree repeater networks (directed and undirected, asymmetric and symmetric) and obtain the analytical expressions of the average of the maximum teleportation fidelities for each of these binary tree networks. We contribute a methodology for the analytical calculation of pathlengths in all considered graph types. Based on these, we have used simple Werner state-based models and are able to identify the parameter ranges for which these networks can show quantum advantage. We also explore the role of maximally entangled states in the network to enhance the quantum advantage. We provide a detailed examination of the large-scale behavior of these networks, obtaining the limiting value of the average maximum teleportation fidelity as the number of nodes, $N$, approaches infinity, same as fractal tree. Our findings reveal that the directed symmetric binary tree represents the most advantageous topology for quantum teleportation within this context. From the context of quantum repeater networks, this work makes a significant advancement in the process of identifying resourceful tree networks for distributed quantum teleportation i.e. teleportation between all possible sources and targets.
keywords:
Average of Maximum Teleportation Fidelity, Binary Trees, Fractal Trees, Werner State, Quantum Network
1 Introduction
@@
In the last decade, we have seen a steady rise in the research and implementation of quantum networks with the vision of creating a full-stack quantum internet [1, 2, 3, 4, 5, 6]. Quantum networking also becomes crucial in distributed quantum computing, where separated quantum computers can work on a similar algorithm [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. Not only computing, but quantum networks can become a test bed for various information processing tasks with many more parties and distributed scenarios. In that spirit, we have seen network topology to play a significant role and value-addition in choosing a network best suited for a particular task [19]. Unlike a classical network [20, 21, 22, 23, 24, 25, 26, 27, 28, 29], one of the unexplored areas is the large-scale quantum networks. For a practical purpose, quantum networks can be deployed either through satellite-based free space technology [30, 31] or large ground-based repeater technology [32, 33, 34, 18]. In an entanglement-based repeater network, entanglement, being the key ingredient, have a crucial role in various communication tasks like sending quantum information, classical information, key generation [35, 36, 37] and secret sharing [38, 39, 40, 41, 42, 43, 44, 45, 36, 46, 47, 48, 49, 50, 51, 52]. In such networks the distribution of entanglement (both bipartite [34, 18] and multipartite [53]) becomes a crucial factor despite limitations of achieving relevant quantum advantages [54, 55, 13]. In this article, we consider a ground-based repeater network specially curated for the purpose of teleportation.
@@
In a repeater-based quantum network scenario, establishment of entanglement between two distant nodes [34, 61, 18] happens through the process of entanglement swapping in the repeater stations [62]. Given the fidelity of intermediate states, it becomes interesting to find out the teleportation fidelity of the entangled state established between the source and target as a result of swapping at the repeater station (Fig. 1). In a quantum repeater network made up of Werner states (parametrized by a set of values $p_{i}$), the maximum teleportation fidelity through a particular path (P) connecting the source (S) and the target (T) can be calculated $F^{\text{tel}}$ as $F^{\rm max}_{{\rm ST},\mathcal{P}}(\rho_{\rm wer})=\left(1+\prod_{i\in\mathcal{P}}p_{i}\right)/2$ [62, 63]. In fact, this is the teleportation fidelity of the final entangled state between S and T created by swapping through that path. In general, in a full-scale quantum network, there can be multiple paths from a source to a target, and hence multiple teleportation fidelities between S and T based on different paths can exist. However, we take the maximum of all the teleportation fidelities when we fix the source and target [19]. To quantify the resourcefulness of a network globally in the context of teleportation, a measure called average of maximum teleportation fidelity ($F^{\text{tel}}_{\text{avg}}$) achieved by all paths was introduced in a recent work [19]. This has given us a global metric ($F^{\text{tel}}_{\text{avg}}$) to estimate the capability of a network in carrying out teleportation over any source and target, and eventually ranking them. In that article [19], the authors have investigated how teleportation fidelity (on an average scale) varies in both loop and loopless networks. We have examined tree sequences that range from chains with all nodes having degree two (apart from the extreme two nodes) to stars, where the hub has the most links. Other trees, like branching trees, that can be created by gradually rearranging nodes from the chain, have also been investigated [19]. It’s been demonstrated that the network can achieve the quantum average fidelity with a lower average distance. In the case of a repeater-based distributed quantum protocol, the intermediate state can get noisier, and in that case, the efficient protocol proposed in [64] can be used to ensure the optimal teleportation fidelity.
@@
Motivated by this newly introduced measure $F_{\text{avg}}^{\text{tel}}$ in undirected simple trees, we focus here on a class of directed and undirected binary trees, a subclass of the Cayley tree. A Cayley tree is an acyclic, connected graph in which each node has up to $k$ children. The binary trees are a specific subcategory of Cayley tree. A left child and a right child are the maximum number of offspring that each node in the binary tree, a hierarchical tree structure, can have. Note that binary trees are used in computer science for designing algorithms and data structures [65, 66]. A lot of studies have been conducted to investigate the role of (un)directed trees on percolation [67, 68, 69, 70, 71, 72], synchronization [73, 74], ecological stability, and in the immune network model [75, 76]. Investigating the average of maximum teleportation fidelity in directed and undirected binary trees is necessary to understand the process of quantum information transfer in a hierarchical network structure and, hence, will be the focus of the current work. Specifically, both symmetric and asymmetric binary trees will be used to study the $F_{\text{avg}}^{\text{tel}}$ in this article. In the symmetric situation, each node has two children (left and right) at each hop/level. The right child will not be taken into consideration if each level has an asymmetric instance. Note that increasing the depth of an undirected binary tree can lead to a fractal tree [77]. This scenario can be mapped to the undirected symmetric binary tree. The current work unfolds two aspects: (i) How does symmetry in a binary graph shape the global teleportation fidelity, and (ii) how does directionality in links influence the global behavior?
@@
In this article, we consider four types of binary trees: directed asymmetric binary tree (DABT), directed symmetric binary tree (DSBT), undirected asymmetric binary tree (UABT) and undirected symmetric binary tree (USBT), in which the USBT will be a fractal when $d$ is very large. For each case, we analytically calculate the number of shortest paths of different pathlengths. Further, based on these numbers, we analytically and numerically find out the average of maximum teleportation fidelities for all these binary tree topologies (See Section 2.3, Fig. 2). Our network models are based on edges made up of Werner states parametrized by the parameter $p$ uniformly over the network. We also obtain the range of $p$ for which each networks show quantum advantage with varying network size $N(d)$, (parametrized by depth $d$). In that process, we also find out the limiting value of the average of the maximum teleportation fidelity for each of these networks for a large number of $N(d)$. Given that each Werner state is useful for teleportation for $p>1/3$, we find the lower bound of the average of the maximum teleportation fidelity for each network and capture its variation with increasing $N(d)$.
@@
In Section 2, we give a brief introduction to existing concepts like entanglement swapping, average of maximum teleportation fidelity and binary tree network. In Section 3, we obtain the average of maximum teleportation fidelity for each of these four binary networks. In Section 4, we analyze the quantum advantage for each of these networks and also find the lower bounds of the average of maximum teleportation fidelities when each link is useful for teleportation. In Section 5, we have calculated the average of maximum fidelity when the input state parameter of the Werner states is different for each edge of the corresponding binary tree network (the input parameters are chosen from a uniform distribution). Finally, in Section 6, we have concluded by summarizing the findings of this work.
@@
Figure 1: Entanglement Swapping: (a) Alice and Charlie are not connected, but entangled pairs are shared between Alice-Bob and Bob-Charlie. (b) Bob’s measurement creates entanglement between Alice and Charlie. (c) Sequential protocol to establish entanglement between $N+2$ nodes using NN swaps.
2 Preliminaries
@@
2.1 Entanglement Swapping and Quantum Repeater
Entanglement swapping is the process by which one can teleport a quantum entanglement from one location to another, thus by connecting two nodes that are not initially entangled with each other. The simplest scheme of entanglement swapping involves three parties, Alice, Bob and Charlie. Here Alice and Charlie shares Bell state $|\Phi^{+}\rangle_{AB_{1}}$ and $|\Phi^{+}\rangle_{B_{2}C}$ with Bob. Now, if Bob carries out two two-qubit Bell-state measurements on his qubits $B_{1}$ and $B_{2}$, Alice and Charlie will end up with two two-qubit states depending upon the measurement outcomes (See Fig. 1) at Bob’s location,
@@ 2.1 Entanglement Swapping and Quantum Repeater
as a resource between Alice-Bob and Bob-Charlie, respectively. After first swapping the state between Alice and Charlie, will become [62, 63],
@@
If we extend this idea to a $(N+2)$ node scenario, to establish the connection between the 11st and the $(N+2)$th node, then there should be $N$ such swappings (See Fig. 1(c)). The resultant state will become,
@@
2.2 Teleportation Fidelity and Average of Maximum Teleportation Fidelity
@@
2.3 Binary Tree Network
A binary tree is one of the hierarchical tree data structures, with each node having at most two child nodes.Scalable and efficient network communication is achieved through the widespread use of hierarchical tree structures in communication networks for routing, multicast distribution, and efficient packet forwarding, all of which are well studied in classical network scenarios [65, 66]. In addition to these, the tree cluster scenarios are used in quantum repeater architecture to enable long-range quantum communication [79], and tree-encoded photonic cluster states are utilised to enable loss-tolerant quantum memory/information storage [80]. Furthermore, the binary tree serves as the basic unit of fractal tree structures, as discussed in the literature, and represents a special case of the Cayley tree. Based on the structure, there are two types of binary trees: asymmetric and symmetric. Similarly, based on the directionality of the edges, we can classify binary trees into two types: directed and undirected. Here in this article, we consider four different types of binary trees: directed asymmetric and symmetric binary trees (DABT, DSBT) and undirected asymmetric and symmetric binary trees (UABT, USBT). All other binary tree configurations can be regarded as intermediate cases lying between these four extremes. These limiting cases also illustrate how the average teleportation fidelity varies with dimensionality and directionality.
@@ 2.3 Binary Tree Network
In Figs. 2(a), 2(b), 2(c), and 2(d), we showed four different cases of binary tree network i.e. directed asymmetric binary tree (DABT), directed symmetric binary tree (DSBT), undirected asymmetric binary tree (UABT), and undirected symmetric binary tree (USBT) respectively. The root node (depicted in skyblue colour) of the tree is node number 1 in each of the cases. In the case of the asymmetric trees (DABT and UABT), the leaf nodes (depicted in green colour) are 3, 5, 6, and 7, whereas the leaf nodes are 8, 9, 10, 11, 12, 13, 141 and 15 in the case of symmetric trees (DSBT and USBT). The depth of each tree is $d=3$ in Figs. 2(a), (b), (c), and (d). The total number of nodes in DABT, DSBT, UABT and USBT is 7, 15, 7, and 15, respectively.
@@
Figure 2: Top Panel: We have (a) DABT and (b) DSBT with 7 and 15 nodes, respectively. Bottom Panel: We have (c) UABT and (d) USBT with 7 and 15 nodes, respectively. The depth of each binary tree is $d$ = 3.
Note that a fractal tree (FT) can be created by splitting one branch into two sub-branches, which is symmetric, and doing this step infinitely. The fractal trees are studied in the context of telecommunications as well [81, 82]. This tree is similar to the case of the undirected symmetric binary tree. As depicted in Fig. 3, if we constantly increase the depth of the undirected symmetric binary tree, then it will lead to a fractal tree when $d$ is very large.
Figure 3: The fractal tree from an undirected symmetric binary tree. From left to right, the depths of the binary trees are 2, 4, 6, 8, and 10, respectively.
3 Average Teleportation Fidelity in Binary Tree Network
In this section, we consider a repeater-based entangled network of $N$ nodes arranged in a binary tree structure. All the edges of our network are made up of Werner states given in Eq. (2) parameterized by the parameter $p$. In this section, we consider four cases of the binary tree network: 1) directed asymmetric binary tree (DABT), 2) directed symmetric binary tree (DSBT), 3) undirected asymmetric binary tree (UABT), 4) undirected symmetric binary tree (USBT) and obtain the analytical expression for average teleportation fidelity for each of these networks. In practical cases, an intermediate link will be destroyed after a swapping process. Therefore, communication between the source and target pair is impossible if there is only one Werner state present between any two nodes. Therefore, there is an ensemble of Werner states between any source-target pair. Due to this fact, the qubit can be transferred from any source and target pair. This is the reason we considered all possible paths in the binary tree network.
Figure 4: Toy Example: In this UABT network with 7 nodes, the first connection is established between 1 and 6 (shown using the red line). After that, the connection is established between 1 and 5 (orange line) and finally between 1 and 7 (green line).
3.1 Notations
The related and useful notation is shown in the Table 1 . We have used the same notations through out the article until stated otherwise.
Notation Explanation
$p$ Werner state parameter
$d$ The depth of the binary tree
NN The number of nodes in the binary tree
$l_{d}^{(i)}$ the number of $i$ length path at level $d$
$\mathcal{F}_{i}$ Teleportation fidelity of $i$ length path
$F^{\text{tel}}_{\text{avg}}$ Average of maximum teleportation fidelity
Table 1: Notations
3.2 Directed Binary Tree
In this subsection, we consider two types of directed binary trees. These include both asymmetric and symmetric cases. Here, the directed link indicates that teleportation is permitted only in one direction. As an example, in Fig. 2(a) and (b), a directed link from node 1 to node 6 implies that teleportation can occur from node 1 to node 6, but not from node 6 to node 1, and so on.
@@ 3.2 Directed Binary Tree
$\left.F^{\text{tel}}_{\text{avg}}(d,p)\right|_{\text{DABT}}=\frac{\sum\limits_{i=1}^{d}[2+2(d-i)]\mathcal{F}_{i}}{\sum\limits_{i=1}^{d}[2+2(d-i)]}=\frac{1}{2}+\frac{p[p(p^{d}-1)+d(1-p)]}{d(d+1)(1-p)^{2}}.$ (7)
@@
$\left.F^{\text{tel}}_{\text{avg}}(N,p)\right|_{\text{DABT}}=\frac{\sum\limits_{i=1}^{(N-1)/2}(N+1-2i)\mathcal{F}_{i}}{\sum\limits_{i=1}^{(N-1)/2}(N+1-2i)}=\frac{1}{2}+\frac{2p[2p(p^{\frac{N-1}{2}}-1)+(N-1)(1-p)]}{(N-1)(N+1)(1-p)^{2}}.$ (8)
@@
$\left.F^{\text{tel}}_{\text{avg}}(d,p)\right|_{\text{DSBT}}=\frac{\sum\limits_{i=1}^{d}[2^{d+1}-2^{i}]\mathcal{F}_{i}}{\sum\limits_{i=1}^{d}[2^{d+1}-2^{i}]}=\begin{cases}\displaystyle\frac{1}{2}+\frac{2^{d+1}-d-2}{2\left[(d-1)\cdot 2^{d+1}+2\right]},&\text{if }~p=\frac{1}{2},\\[12.0pt] \displaystyle\frac{1}{2}+\frac{p\left[(2^{d}-1)-p(2^{d+1}-1)+2^{d}p^{d+1}\right]}{2(1-p)(1-2p)\left[(d-1)\cdot 2^{d}+1\right]},&\text{otherwise},\end{cases}$ (9)
@@
$\left.F^{\text{tel}}_{\text{avg}}(N,p)\right|_{\text{DSBT}}$ $=$ $\frac{\sum\limits_{i=1}^{\log_{2}[(N+1)/2]}(N+1-2^{i})\mathcal{F}_{i}}{\sum\limits_{i=1}^{\log_{2}[(N+1)/2]}(N+1-2^{i})}$ (10)
$=$ $\begin{cases}\displaystyle\frac{1}{2}+\frac{N-\log_{2}(N+1)}{2\left[(N+1)\log_{2}(N+1)-2N\right]},&\text{if }~p=\frac{1}{2},\\[12.0pt] \displaystyle\frac{1}{2}+\frac{p[N(1-2p)+(N+1)p^{\log_{2}(N+1)}-1]}{4(1-p)(1-2p)\left[(N+1)\log_{2}(N+1)-2N\right]},&\text{otherwise}.\end{cases}$ (11)
A detailed derivation of the formula is provided in appendix A.1. For both the cases of directed binary trees, $p=1$ leads to $F^{\text{tel}}_{\text{avg}}(N,p)=1$.
3.3 Undirected Binary Trees
In this subsection, we consider two types of undirected binary trees, which incorporate both asymmetric and symmetric cases. Here, the undirected link indicates bidirectional teleportation. As an example, in Figs. 2(c) and 2(d), teleportation can occur from node 1 to node 6 as well as from node 6 to node 1, and so on i.e. no restriction due to the directionality.
Undirected Asymmetric Binary Trees (UABT) : Here we consider an undirected asymmetric binary tree network (UABT) where each link represents a Werner state, where each Werner state is parameterized by the classical mixing parameter $p$. It is shown in Fig. 2(c). In the UABT network, we start with $3$ nodes at level $h=1$, and after that we add two nodes at every level. So in that case, $N$ will be odd numbers starting from $3$ (i.e. $N=3,5,7\cdots$).